A field is 3 metres longer than wide.
The width of the field is x metres.
The area of the field is 10m squared.
Find x

Found 2 solutions by fractalier, addingup:
Answer by fractalier(6550)

If the width is x, ten the length is x+3.
Area A = LW = x(x+3) = 10
x^2 + 3x = 10
x^2 + 3x - 10 = 0
(x+5)(x-2) = 0
x = 2
It is a 2m by 5m field.

Answer by addingup(3677)

L= W+3 and we are calling the Width x:
L= x+3
And we are told that:
x*(x+3)= 10
x^2+3x= 10 Subtract 10 from both sides:
x^2+3x-10= 0 Factor the equation (FOIL). Find two numbers that when added or subtracted give you 3 and when multiplied -10. The factors of 10 are: 1, 2, 5, 10. Got them, 5 and 2 are our numbers:
(x+5)(x-2)= 0 Now break in 2 equations:
x+5= 0 or x-2= 0
x= -5 or x= 2 We are looking for a positive number, so toss the -5 and keep the 2:
x= 2
Proof:
x*(x+3)= 10
2*(2+3)= 10
2*5= 10 We've got the correct answer.